# CLIP Meets Diffusion: A Synergistic Approach to Anomaly Detection

### v1

@@ 4.2 Anomaly Classification
This means that the score becomes larger when the degree of focus of the cross-attention map is high, that is, when the maximum value is significantly greater than the median value.
In the few-shot setup, we additionally aggregate the vision-guided score maps, defining the scores as follows in accordance with common practice [25]:
@@ 5 Comparisons
We use the MVTec-AD [2] and VisA [60] datasets, which are mainly used as benchmark datasets in anomaly detection. Each dataset consists of several object categories, and each object category is divided into a training set containing only normal images and a test set containing a mixture of normal and abnormal images. The abnormal images have various types of anomalies. Query images are from the test set, and reference images are sampled from the training set.
@@
| 0-shot | CLIP-AC [37] | 74.00.0 | 89.10.0 | 59.30.0 | 67.00.0 |
| WinCLIP [25] | 91.80.0 | 96.50.0 | 78.10.0 | 81.20.0 | |
@@
| 1-shot | PaDiM [11] | 76.63.1 | 88.11.7 | 62.85.4 | 68.34.0 |
| PatchCore [43] | 83.43.0 | 92.21.5 | 79.92.9 | 82.82.3 | |
| WinCLIP [25] | 93.12.0 | 96.50.9 | 83.84.0 | 85.14.0 | |
| PromptAD [29] | 92.61.0 | 95.70.9 | 83.12.1 | 85.71.2 | |
@@
| 2-shot | PaDiM [11] | 78.93.1 | 89.31.7 | 67.45.1 | 71.63.8 |
| PatchCore [43] | 86.33.3 | 93.81.7 | 81.64.0 | 84.83.2 | |
| WinCLIP [25] | 94.41.3 | 97.00.7 | 84.62.4 | 85.82.7 | |
| PromptAD [29] | 93.21.0 | 96.11.0 | 85.01.8 | 87.21.8 | |
@@
| 4-shot | PaDiM [11] | 80.42.5 | 90.51.6 | 72.82.9 | 75.62.2 |
| PatchCore [43] | 88.82.6 | 94.51.5 | 85.32.1 | 87.52.1 | |
| WinCLIP [25] | 95.21.3 | 97.30.6 | 87.31.8 | 88.81.8 | |
| PromptAD [29] | 93.80.7 | 96.50.7 | 86.21.6 | 88.11.9 | |
@@
Specifically, we compare our method, CLIPFUSION, with methods utilizing vision models (PaDiM, PatchCore) and those utilizing vision-language models (CLIP-AC, Trans-MM, MaskCLIP, WinCLIP, PromptAD333The authors of PromptAD use the test set as a validation set during training to record the best performance for each category separately and report the aggregated results. For a fair comparison, we treat the training epoch as a fixed hyperparameter across categories and report the best-performing checkpoint reproduced using the official codebase. See Appendix F for more details.). Note that, among these methods, only CLIPFUSION and WinCLIP are fully capable of handling both segmentation and classification tasks in both zero-shot and few-shot scenarios.
Experimental details
5.1 Anomaly Segmentation
Table 1 presents a comparison of our method’s performance with baseline methods under zero-shot and few-shot setups for anomaly segmentation. The results show that CLIPFUSION consistently outperforms the baseline methods across all setups, datasets, and metrics. Our method, in particular, demonstrates dramatic performance improvements in the zero-shot setup. This highlights the critical contribution of the diffusion-based model’s cross-attention map to zero-shot anomaly segmentation. Additionally, the significant performance improvements in the few-shot setup demonstrate how the diffusion-based model complements the CLIP-based model.
@@ 6 Empirical Study
In this section, all results are reported under the one-shot setup for consistency.
@@
6.1 Ablation Study
Table 3 presents the comparative performance of different models. As expected, the diffusion-based model, which captures local features effectively, performs better in the segmentation task, while the CLIP-based model, which captures global features effectively, performs better in the classification task. The combination of different models consistently yields better results compared to using either model alone. This empirical evidence suggests that the two models capture complementary features, making their integration highly beneficial for both segmentation and classification tasks.
Table 4 presents the comparative performance of different modalities. The results highlight the complementary strengths of both modalities. Combining the two modalities enhance performance across tasks by leveraging both semantic and spatial information.
@@ 6.2 Extraction Strategies from Diffusion Denoisers
Table 5 highlights the impact of matching timestep and block numbers when calculating (Equation 6). Matching higher timesteps with lower blocks yields better performance, and vice versa (Timestep 201 → Block 3, Timestep 401 → Block 2, Timestep 801 → Block 1). This pattern reflects the intrinsic behavior of the diffusion denoiser: at higher timesteps, the denoiser interprets the input image as highly noisy and prioritizes the extraction of global features. The lower blocks, which process coarse-grained information, are particularly well-suited for this task. Using higher timesteps can be seen as a regularization strategy, as they suppress local details and highlight broader patterns.
@@ 6.3 One-for-all paradigm
IIPAD [34] recently proposes a new setting for few-shot anomaly detection, termed the one-for-all paradigm. It aims to train a single unified model that can detect anomalies across multiple object categories. In this paradigm, the one-shot setting refers to using normal images (one for each category), where is the number of categories. Table 7 presents a comparison with IIPAD (the SOTA method under this paradigm) on the MVTec-AD dataset in the one-shot setting, where our method achieves higher performance. See Appendix G for more results.
6.4 Inference speed
| Latency [ms] | 0-shot | 1-shot | 2-shot | 4-shot |
| WinCLIP | 4782 | 9514 | 9633 | 9710 |
| CLIPFUSION | 783 | 1240 | 1269 | 1278 |
Our method can be viewed as an extension of WinCLIP. We report the inference latency measured on an NVIDIA A100 GPU in Table 8, with a batch size of 1. As shown, our method is approximately 6–8 times faster than WinCLIP. To capture fine-grained details, WinCLIP slides multi-scale windows over the query image, resulting in hundreds of forward passes. In contrast, our CLIP-based model utilizes patch embeddings obtained from a single forward pass. On top of this, we employ a diffusion-based model not as a denoiser for iterative generation, but as a feature extractor to capture fine-grained details. As a result, our method requires only a few forward passes in total.
@@ 7 Conclusion
We propose a novel methodology that synergistically combines discriminative and generative foundation models, leveraging their complementary strengths. As a concrete application, we introduce CLIPFUSION, which integrates CLIP and diffusion models for anomaly detection and consistently outperforms baselines in segmentation and classification across zero-shot and few-shot scenarios. The holistic nature of this approach not only advances anomaly detection but also demonstrates the potential of multi-modal and multi-model integration for complex recognition tasks.
Appendix A Further preliminaries on diffusion models
A diffusion model [21] is trained through a sequential process that involves forward diffusion and backward denoising. In the diffusion process, random noise is progressively added to a clean image over several steps, following the equation:
where is the noise schedule hyperparameter, which determines the amount of noise added to the image [36]. In the denoising process, random noise is progressively removed over several steps using the following equation to restore the original image:
In this process, the denoiser learns to predict the added noise from the noisy image at a given time step [49]. During inference, random noise is sampled, and the denoiser iteratively removes it to generate a new image. By conditioning on a text prompt , the denoiser guides the generation process to align the resulting image with the provided text. Diffusion models demonstrate superior performance in understanding image details by generating high-quality images compared to other generative models [13].
Appendix B Further motivation for feature-based anomaly detection
@@
When utilizing generative models in anomaly detection, the reconstruction-based approach has traditionally been the dominant paradigm [53, 54, 18]. This approach compares the original image and the reconstructed image at both ends. However, since generative models are primarily designed for creative image generation, they can produce high reconstruction errors even for normal images. In addition, deciding how to apply the mask, including its location and size is challenging.444Since the location of the anomalous part is not known, it is unclear where to apply the mask. Also, if the mask is too large, the reconstructed image may deviate significantly from the original image. Conversely, if the mask is too small and fails to fully cover the anomaly, the anomaly may persist or even extend in the reconstructed image. As a result, the feature-based approach has recently gained prominence.
@@
Appendix C Further experimental details
@@
For PatchCLIP, we use the same pre-processing pipeline as in WinCLIP [25], which is specified in OpenCLIP [24], for both MVTec-AD and VisA datasets. First, bicubic interpolation is used to resize images to a resolution of 240 to fit the ViT-B-16-plus-240 model. After which, channel-wise standardization is applied with the precomputed values of (0.48145466, 0.4578275, 0.40821073) for the mean and (0.26862954, 0.26130258, 0.27577711) for the standard deviation. For MapDiff, we use bilinear interpolation to resize images to a resolution of 512 before feeding into a customized Stable Diffusion Inpainting pipeline from the diffusers library [50].
@@
The prompts are derived from those used to train CLIP for the ImageNet dataset [12]. For example, in the case of a segmentation task on MVTec-AD, PatchCLIP uses the prompt a good, cropped picture of the [state] [object] for classification," where [state] is either “damaged" or “perfect." MapDiff uses the prompt a close-up cropped png photo of a [object] with [state] for anomaly segmentation," where [state] includes “crack," “hole," “residue," and “damage," for query images, and “a photo of a perfect [object]" for reference images.
@@
We follow previous studies [10, 11, 43, 25] and use the following performance metrics. For anomaly classification, we report the Area Under the ROC Curve (AUROC) at the image level and the Area Under the Precision-Recall Curve (AUPR). For anomaly segmentation, we report the AUROC at the pixel level and the Area Under the Per-Region Overlap Curve (AUPRO) [3, 28]. Both AUPR and AUPRO are particularly sensitive to class imbalance.
@@
The performance of Trans-MM, MaskCLIP, PaDiM, PatchCore, CLIP-AC, and WinCLIP are referenced from the WinCLIP paper [25].
Appendix D Further empirical study
@@
Table 10 illustrates the impact of using state ensemble for a cross-attention map. Without ensemble, only a generic term “damage" is used to generate the cross-attention map. In contrast, the ensemble approach averages cross-attention maps derived from both the generic term “damage" and more specific terms such as “crack," “hole," and “residue." This approach significantly improves performance across tasks. By incorporating these multiple specific states, the method generates diverse cross-attention maps that capture fine-grained and complementary aspects of anomalies.
@@
| Cable | “crack", “hole", “residue", “damage" | 94.6 |
| “crack", “poke", “scratch" | 95.0 | |
| Pill | “crack", “hole", “residue", “damage" | 83.0 |
| “crack", “scratch", “residue" | 83.5 |
Appendix E Limitations
Our method leverages pretrained foundation models (e.g., CLIP and Stable Diffusion) without any task-specific fine-tuning. While this reduces adaptability to domain-specific anomalies, the method still generalizes well in zero-shot and few-shot scenarios across diverse object categories. Investigating how fine-tuning these foundation models for specific use cases could further improve performance can be an interesting direction for future research.
Appendix F Limitations of few-shot training
Training-based methods inherently suffer from epoch selection sensitivity. Figure 4 shows that the epoch at which PromptAD performs best differs significantly across object categories. In conventional training setups, such epoch selection is guided by a validation set. However, in few-shot anomaly detection, where only one to four images are available, it is practically hard to reserve any for a separate validation set. The authors of PromptAD circumvent this by using the test set for validation in their official implementation.555https://github.com/FuNz-0/PromptAD However, test sets are not accessible during model training in real-world applications. In contrast, CLIPFUSION is training-free and uses few-shot examples solely for feature extraction, making validation sets and epoch selection unnecessary.
Figure 5 reveals that even when the training epoch is fixed, performance varies considerably across different random seeds. This highlights the instability and seed sensitivity of few-shot training, making it difficult to ensure consistent and reproducible results without a validation set.
Appendix G Further results in the one-for-all paradigm
We evaluate anomaly detection performance under the one-for-all paradigm, where a single model is applied across all object categories. Recall that, in this paradigm, the one-shot setting provides one image per category, for a total of images across object categories. IIPAD trains a prompt generator using these examples for the one-for-all setting. In contrast, CLIPFUSION remains entirely training-free. Nevertheless, it outperforms or matches IIPAD across all configurations and datasets for both segmentation and classification.

### v3

@@ 4.2 Anomaly Classification
This means that the score becomes larger when the degree of focus of the cross-attention map is high, that is, when the maximum is significantly greater than the median.
In the few-shot setup, we additionally aggregate the vision-guided score maps, defining the scores as follows in accordance with common practice (jeong2023winclip):
@@ 5 Comparisons
We use the MVTec-AD (bergmann2019mvtec) and VisA (zou2022spot) datasets, which are mainly used as benchmark datasets in anomaly detection. Each dataset consists of several object categories, and each object category is divided into a training set containing only normal images and a test set containing a mixture of normal and abnormal images. Query images are from the test set, and reference images are sampled from the training set.
@@
Specifically, we compare our method, CLIPFUSION, with methods utilizing vision models (PaDiM, PatchCore) and those utilizing vision-language models (CLIP-AC, Trans-MM, MaskCLIP, WinCLIP, PromptAD333The authors of PromptAD use the test set as a validation set during training to record the best performance for each category separately and report the aggregated results. For a fair comparison, we treat the training epoch as a fixed hyperparameter across categories and report the best-performing checkpoint reproduced using the official codebase. See Appendix A.5 for more details.). Note that, among these methods, only CLIPFUSION and WinCLIP are fully capable of handling both segmentation and classification tasks in both zero-shot and few-shot scenarios.
Experimental details
For PatchCLIP, we employ OpenCLIP (ilharco_gabriel_2021_5143773; cherti2023reproducible; schuhmann2022laionb), which has been pretrained on LAION-400M (schuhmann2021laion), using a Vision Transformer (ViT) (dosovitskiy2020image) as its image encoder. For MapDiff, we utilize the pretrained Stable Diffusion v2 model (Rombach_2022_CVPR), where the denoiser is based on U-Net (ronneberger2015u).
We set in Equation 4 for segmentation and in Equation 9 for classification to reflect the different importance of the models in each task. We report the average performance and standard deviations over five runs with different seeds. Further details are in Appendix A.2.
| Setup | Method | MVTec-AD | VisA | ||
| AUROC | AUPRO | AUROC | AUPRO | ||
| 0-shot | Trans-MM (chefer2021generic) | 57.50.0 | 21.90.0 | 49.40.0 | 10.20.0 |
| MaskCLIP (zhou2022extract) | 63.70.0 | 40.50.0 | 60.90.0 | 27.30.0 | |
| WinCLIP (jeong2023winclip) | 85.10.0 | 64.60.0 | 79.60.0 | 56.80.0 | |
| CLIPFUSION (ours) | 90.90.0 | 87.10.0 | 92.10.0 | 83.30.0 | |
| 1-shot | PaDiM (defard2021padim) | 89.30.9 | 73.32.0 | 89.90.8 | 64.32.4 |
| PatchCore (roth2022towards) | 92.01.0 | 79.72.0 | 95.40.6 | 80.52.5 | |
| WinCLIP (jeong2023winclip) | 95.20.5 | 87.11.2 | 96.40.4 | 85.12.1 | |
| PromptAD (li2024promptad) | 94.90.8 | 87.71.5 | 95.70.9 | 84.01.2 | |
| CLIPFUSION (ours) | 95.80.1 | 90.40.1 | 97.30.1 | 87.00.3 | |
| 2-shot | PaDiM (defard2021padim) | 91.30.7 | 78.21.8 | 92.00.7 | 70.12.6 |
| PatchCore (roth2022towards) | 93.30.6 | 82.31.3 | 96.10.5 | 82.62.3 | |
| WinCLIP (jeong2023winclip) | 96.00.3 | 88.40.9 | 96.80.3 | 86.21.4 | |
| PromptAD (li2024promptad) | 95.00.7 | 88.31.0 | 96.30.9 | 84.80.8 | |
| CLIPFUSION (ours) | 96.30.1 | 91.40.2 | 97.70.1 | 87.30.2 | |
| 4-shot | PaDiM (defard2021padim) | 92.60.7 | 81.31.9 | 93.20.5 | 72.61.9 |
| PatchCore (roth2022towards) | 94.30.5 | 84.31.6 | 96.80.3 | 84.91.4 | |
| WinCLIP (jeong2023winclip) | 96.20.3 | 89.00.8 | 97.20.2 | 87.60.9 | |
| PromptAD (li2024promptad) | 95.50.6 | 89.00.6 | 96.30.8 | 85.10.9 | |
| CLIPFUSION (ours) | 96.80.1 | 92.10.2 | 98.00.1 | 87.70.5 | |
5.1 Anomaly Segmentation
Table 1 presents a comparison of our method’s performance with baseline methods for anomaly segmentation. The results show that CLIPFUSION consistently outperforms the baseline methods across all setups, datasets, and metrics. Our method, in particular, demonstrates dramatic performance improvements in the zero-shot setup. This highlights the critical contribution of the diffusion-based model’s cross-attention map to zero-shot anomaly segmentation. Additionally, the significant performance improvements in the few-shot setup demonstrate how the diffusion-based model complements the CLIP-based model.
@@ 5.1 Anomaly Segmentation
| 0-shot | CLIP-AC (radford2021learning) | 74.00.0 | 89.10.0 | 59.30.0 | 67.00.0 |
| WinCLIP (jeong2023winclip) | 91.80.0 | 96.50.0 | 78.10.0 | 81.20.0 | |
@@
| 1-shot | PaDiM (defard2021padim) | 76.63.1 | 88.11.7 | 62.85.4 | 68.34.0 |
| PatchCore (roth2022towards) | 83.43.0 | 92.21.5 | 79.92.9 | 82.82.3 | |
| WinCLIP (jeong2023winclip) | 93.12.0 | 96.50.9 | 83.84.0 | 85.14.0 | |
| PromptAD (li2024promptad) | 92.61.0 | 95.70.9 | 83.12.1 | 85.71.2 | |
@@
| 2-shot | PaDiM (defard2021padim) | 78.93.1 | 89.31.7 | 67.45.1 | 71.63.8 |
| PatchCore (roth2022towards) | 86.33.3 | 93.81.7 | 81.64.0 | 84.83.2 | |
| WinCLIP (jeong2023winclip) | 94.41.3 | 97.00.7 | 84.62.4 | 85.82.7 | |
| PromptAD (li2024promptad) | 93.21.0 | 96.11.0 | 85.01.8 | 87.21.8 | |
@@
| 4-shot | PaDiM (defard2021padim) | 80.42.5 | 90.51.6 | 72.82.9 | 75.62.2 |
| PatchCore (roth2022towards) | 88.82.6 | 94.51.5 | 85.32.1 | 87.52.1 | |
| WinCLIP (jeong2023winclip) | 95.21.3 | 97.30.6 | 87.31.8 | 88.81.8 | |
| PromptAD (li2024promptad) | 93.80.7 | 96.50.7 | 86.21.6 | 88.11.9 | |
@@ 6 Empirical Study
In this section, results are reported under the one-shot setup.
6.1 Ablation Study
Table 6 presents the comparative performance of different models. As expected, the diffusion-based model, which captures local features effectively, performs better in the segmentation task, while the CLIP-based model, which captures global features effectively, performs better in the classification task. The combination of different models consistently yields better results compared to using either model alone. This empirical evidence suggests that the two models capture complementary features, making their integration beneficial for both segmentation and classification tasks.
@@ 6.1 Ablation Study
Table 6 presents the performance of different modalities, highlighting their complementary strengths through the combination of semantic and spatial information.
@@ 6.2 Extraction Strategies from Diffusion Denoisers
Table 6 highlights the impact of matching timestep and block numbers when calculating (Equation 8). Note that the decoder consists of four blocks (Block 0 to Block 3). Matching higher timesteps with lower blocks yields better performance, and vice versa (Timestep 201 → Block 3, Timestep 401 → Block 2, Timestep 801 → Block 1). This pattern reflects the intrinsic behavior of the diffusion denoiser: at higher timesteps, the denoiser interprets the input image as highly noisy and prioritizes the extraction of global features. The lower blocks, which process coarse-grained information, are particularly well-suited for this task. Using higher timesteps can be seen as a regularization strategy, as they suppress local details and highlight broader patterns.
@@
6.3 Inference speed
Our method can be viewed as an extension of WinCLIP (jeong2023winclip). We report the inference latency measured on an NVIDIA A100 GPU (40 GB) in Table 7, with a batch size of 1. As shown, our method is approximately 6–8 times faster than WinCLIP.
| Latency [ms] | 0-shot | 1-shot | 2-shot | 4-shot |
| WinCLIP | 4782 | 9514 | 9633 | 9710 |
| CLIPFUSION | 783 | 1240 | 1269 | 1278 |
To capture fine-grained details, WinCLIP slides multi-scale windows over the query image, resulting in hundreds of forward passes. In contrast, our CLIP-based model utilizes patch embeddings obtained from a single forward pass. On top of this, we employ a diffusion-based model not as a denoiser for iterative generation, but as a feature extractor to capture fine-grained details. As a result, our method requires only a few forward passes in total.
@@ 6.4 One-for-all paradigm
IIPAD (lvone) recently proposes a new setting for few-shot anomaly detection, termed the one-for-all paradigm. It aims to train a single unified model that can detect anomalies across multiple object categories. In this paradigm, the one-shot setting refers to using normal images (one for each category), where is the number of categories. Apart from this, we follow the same experimental setting as described in Section 5. Table 8 presents a comparison with IIPAD (the SOTA method under this paradigm) on the MVTec-AD dataset in the one-shot setting, where our method achieves higher performance. See Appendix A.4 for more results.
@@ 7 Conclusion
We propose a methodology that synergistically combines discriminative and generative foundation models, leveraging their complementary strengths. As a concrete application, we introduce CLIPFUSION, which integrates CLIP and diffusion models for anomaly detection and consistently outperforms baselines in segmentation and classification across zero-shot and few-shot scenarios. This holistic use of features from discriminative and generative models may benefit the research community in tackling complex recognition tasks.
Acknowledgments
This work was conducted by Center for Applied Research in Artificial Intelligence(CARAI) grant funded by Defense Acquisition Program Administration(DAPA) and Agency for Defense Development(ADD) (UD230017TD).
Appendix A Appendix
A.1 Further motivation for feature-based anomaly detection
@@
When utilizing generative models in anomaly detection, the reconstruction-based approach has traditionally been the dominant paradigm (zavrtanik2021draem; zhang2023unsupervised; guo2023template). This approach compares the original image and the reconstructed image at both ends. However, since generative models are primarily designed for creative image generation, they can produce high reconstruction errors even for normal images. In addition, deciding how to apply the mask, including its location and size is challenging.444Since the location of the anomalous part is not known, it is unclear where to apply the mask. Also, if the mask is too large, the reconstructed image may deviate significantly from the original image. Conversely, if the mask is too small and fails to fully cover the anomaly, the anomaly may persist or even extend in the reconstructed image. As a result, the feature-based approach has recently gained prominence.
@@
A.2 Further experimental details
@@
For PatchCLIP, we use the same pre-processing pipeline as in WinCLIP (jeong2023winclip), which is specified in OpenCLIP (ilharco_gabriel_2021_5143773), for both MVTec-AD and VisA datasets. First, bicubic interpolation is used to resize images to a resolution of 240 to fit the ViT-B-16-plus-240 model. After which, channel-wise standardization is applied with the precomputed values of (0.48145466, 0.4578275, 0.40821073) for the mean and (0.26862954, 0.26130258, 0.27577711) for the standard deviation. For MapDiff, we use bilinear interpolation to resize images to a resolution of 512 before feeding into a customized Stable Diffusion Inpainting pipeline from the diffusers library (von-platen-etal-2022-diffusers).
@@
The prompts are derived from those used to train CLIP for the ImageNet dataset (deng2009imagenet). For example, in the case of a segmentation task on MVTec-AD, PatchCLIP uses the prompt “a good, cropped picture of the [state] [object] for classification”, where [state] is either “damaged” or “perfect”. MapDiff uses the prompt “a close-up cropped png photo of a [object] with [state] for anomaly segmentation” for query images. The [state] includes commonly observed anomaly types “crack”, “hole”, and “residue”, along with a generic descriptor “damage”. For reference images, it uses the prompt “a photo of a perfect [object]”.
@@
We follow previous studies (cohen2020sub; defard2021padim; roth2022towards; jeong2023winclip) and use the following performance metrics. For anomaly classification, we report the Area Under the ROC Curve (AUROC) at the image level and the Area Under the Precision-Recall Curve (AUPR). For anomaly segmentation, we report the AUROC at the pixel level and the Area Under the Per-Region Overlap Curve (AUPRO) (bergmann2020uninformed; li2021cutpaste). Both AUPR and AUPRO are particularly sensitive to class imbalance.
Hyperparameters
We compare weights of (1, 0), (0, 1), (0.75, 0.25), (0.25, 0.75), and (0.5, 0.5) between the CLIP and diffusion models. We observe a consistent trend: (0.25, 0.75) is effective for segmentation tasks, while (0.75, 0.25) performs better for classification tasks. The same weights are used across all datasets for each task. Please refer to Table 6.
@@
The performance of Trans-MM, MaskCLIP, PaDiM, PatchCore, CLIP-AC, and WinCLIP are referenced from the WinCLIP paper (jeong2023winclip).
A.3 Further empirical study
@@
Table 10 illustrates the impact of using state ensemble for a cross-attention map. Without ensemble, only a generic term “damage” is used to generate the cross-attention map. In contrast, the ensemble approach averages cross-attention maps derived from both the generic term “damage” and more specific terms such as “crack," “hole," and “residue." This approach significantly improves performance across tasks. By incorporating these multiple specific states, the method generates diverse cross-attention maps that capture fine-grained and complementary aspects of anomalies.
@@
| Cable | “crack”, “hole”, “residue”, “damage” | 94.6 |
| “crack”, “poke”, “scratch” | 95.0 | |
| Pill | “crack”, “hole”, “residue”, “damage” | 83.0 |
| “crack”, “scratch”, “residue” | 83.5 |
Table 12 presents results obtained with different sets of states. The results show our method remains reasonably robust across these variations.
| States | Task | Dataset | |
| MVTec-AD | VisA | ||
| “contamination”, “bend”, “cut”, “damage” | Seg. | 95.8 | 97.0 |
| Cls. | 95.4 | 85.9 | |
| “break”, “fold”, “stain”, “damage” | Seg. | 95.8 | 97.3 |
| Cls. | 95.6 | 86.1 | |
A.4 Further results in the one-for-all paradigm
We evaluate anomaly detection performance under the one-for-all paradigm, where a single model is applied across all object categories. Recall that, in this paradigm, the one-shot setting provides one image per category, for a total of images across object categories. IIPAD (lvone) trains a prompt generator using these examples for the one-for-all setting. In contrast, CLIPFUSION requires no additional training. Nevertheless, it outperforms or matches IIPAD across all configurations and datasets for both segmentation and classification, as shown in Tables 13 and 14.